Quantitative Problem: International Machinery Company (IMC) is a Swedish multinational manufacturing company. Currently, IMC's financial planners are considering undertaking a 1-year project in the United States. The project's expected dollar-denominated cash flows consist of an initial investment of $2,900 and a cash inflow the following year of $3,700. IMC estimates that its risk-adjusted cost of capital is 18%. Currently, 1 U.S. dollar will buy 6.0 Swedish kronas. In addition, 1-year risk-free securities in the United States are yielding 2%, while similar securities in Sweden are yielding 1%.

a. If the interest parity holds, what is the forward exchange rate of Swedish krona per U.S. dollar? Do not round intermediate calculations. Round your answer to two decimal places.

   Swedish krona per U.S. dollar

b. If IMC undertakes the project, what is the net present value and rate of return of the project for IMC in home currency? Do not round intermediate calculations. Round your answers to two decimal places.

NPV:    Swedish kronas

Rate of return:    %
Answer #1

Part (a)

Forward exchange rate = Spot rate x (1 + iSweden) / (1 + iUS) = 6 x (1 + 1%) / (1 + 2%) = 5.94 Swedish Kronas per U.S. dollar

Part (b)

Year Linkage 0 1
Cash flows in $ A     (2,900.00)     3,700.00
Exchange rate B              6.00             5.94
Cash flows in Swedish Kronas C = A x B (17,400.00) 21,982.35

NPV = -17,400 + 21,982.35 / (1 + 18%) =  1,229.11 Swedish Kronas

Rate of return = 21,982.35 / 17,400 - 1 = 26.34%